Calculate the maximum possible ATP yields when glucose is completely catabolized to six molecules of CO2 during aerobic respiration.

Explanation please !
Answer #1

The maximum possible ATP yields are 38. They are formed as follows-

1.)Glycolysis - forms 2 molecules of pyruvate, 2ATP and 2NADH.

1 NADH= 3 ATP hence 2NADH = 6 ATP

Total formed in glycolysis = 8 ATP

2.) Link reaction- forms 1 NADH and release one molecule of carbon dioxide

1 NADH= 3 ATP

We should count the link reaction two times because one link reaction uses one pyruvate molecule but one glucose forms two pyruvate molecules.

3.) Kreb's cycle- forms 1ATP, 3NADH and 1 FADH

1 NADH = 3 ATP, 1 FADH= 2 ATP

We should double the kreb's cycle as we did the link reaction.

So total ATP production= ATP from glycolysis +2×( ATP from one link reaction) +2×(ATP from one kreb's cycle)

ATP from kreb's cycle= 1ATP+3×3ATP+1×2ATP=12 ATP

Total ATP production= 8ATP+2×(3ATP) +2×(12ATP)

=8+6+24=38ATP